Derek plans to buy a $26,324.00 car. The dealership offers zero percent financing for 57.00 months with the first payment due at signing (today). Derek would be willing to pay for the car in full today if the dealership offers him $____ cash back. He can borrow money from his bank at an interest rate of 4.96%. SHOW FINANCE CALCULATOR KEY STROKES
Answer #1

Using financial calculator

Set mode to BGN

I/Y=4.96%/12
N=57
PMT=-26324/57
FV=0
CPT PV=$23,506.75

Hence, cashback required=26324-23506.75=$2,817.25